# “How Big is Big Enough?” Adjusting Model Size in Continual Gaussian Processes

### v1

@@ 5 Experiments
Figure 2: (a) Location selection method “greedy variance” with fixed memory (blue curves with $M=10,20,30$ total inducing points per batch) and VIPS, which adaptively selects $M$ (shown at the top) inducing points per batch. (b) (log) Time taken to train the online GP model on the “naval” dataset divided into 20 batches using “greedy variance” selection method with fixed inducing points (oracle: $M=100$, heuristic: $M=1000$) and VIPS, our adaptive method.
5.1 Synthetic data
In the experiment, we aim to illustrate the impact of data distribution on the model’s size requirement during continual learning. In Figure 1, we consider three different datasets, each divided into 10 batches, and record the total number of inducing points selected at each batch by our inducing point method, VIPS. In the first scenario, the model encounters a growing input space, leading to a linear increase in the model’s capacity with each batch. In the second, the data distribution remains consistent across batches. After initial learning, the number of inducing points asymptotes to the constant value $M=30$. In the third scenario, the data is concentrated in a part of the input space with occasional outliers in the last batches. Using a large model from the beginning would be a waste of resources. In contrast, our method maintains a low model size for most batches and only increases dynamically when encountering outliers. These scenarios show VIPS’ ability to manage model size across different data distributions during continual learning. See Appendix F.1 for details.
5.2 Speed and accuracy
In continual learning scenarios, manually setting the appropriate number of inducing points beforehand is not possible since this depends on dataset properties and total data points. Fixed-memory approaches can lead to suboptimal performance if the inducing point set is too small, or computational waste if the model is too large. The following two experiments demonstrate these limitations. In contrast, our proposed method VIPS only increases the number of inducing points as needed, highlighting the importance of an adaptive number of inducing points in continual learning.
Synthetic data. We compare adaptive and fixed memory methods using a synthetic dataset divided into four equal batches. We evaluate the "greedy variance" location selection method [3] under two scenarios: with a fixed number of inducing points and using our stopping criterion to determine the number of inducing points. We consider three fixed settings ($M=10$, $20$, $30$ total number of inducing points per batch) and record the root mean square error (RMSE) on a test set, using an exact GP model with access to all current training data as a benchmark. Figure 2(a) shows that fixed memory methods with 10 and 20 inducing points become less accurate as more data is observed, while our method VIPS maintains accuracy at each batch. Both the fixed memory method with $M=30$ and VIPS match the performance of an exact GP. This demonstrates the importance of model size in continual learning and the need for adaptive methods to determine the number of inducing points.
Naval dataset. We compare the time complexity of the location selection method “greedy variance” [3] under three scenarios. In the first scenario, we assume we have access to an oracle to determine the required number of inducing points for near-exact performance, and we set $M=100$ based on the accuracy obtained on a sparse model compared to an exact GP. In the second scenario, we take a heuristic with a large memory ($M=1000$, around 1/10th of the total data points) to ensure sufficient capacity. In the third scenario, we use our method VIPS to determine the number of inducing points. We use the “naval” UCI dataset divided into 20 batches and record the training time per batch. Figure 2(b) shows that both fixed-memory approaches have higher computational time. In contrast, our VIPS method adaptively adjusts $M$ as needed, leading to a reduction in training time.
Overall these simple experiments aim to demonstrate the advantages of adaptive strategies over fixed approaches, achieving high-quality approximation without wasting computational resources.
5.3 UCI datasets
We compare our inducing point selection method, VIPS, to two other adaptive approaches that iteratively choose the number of inducing points with different stopping criteria: Conditional Variance [3] and OIPS [14]. Technical details of these methods are provided in Appendix E.
We evaluate five datasets from the UCI repository [8], simulating a streaming scenario by sorting the data based on the first dimension and dividing it into batches. Results are reported over five 80%-training and 20%-testing random splits, including root mean square error (RMSE) and negative log predictive density (NLPD). For Conditional Variance, we set $\eta=0.1$, a threshold that consistently achieves good accuracy across different datasets, and for OIPS, we use default parameters, $\rho=0.8$. For VIPS, we set $\delta=0.05$. Table 1 summarises the last batch results for each dataset in terms of inducing points and RMSE, with an exact GP as a benchmark. Conditional Variance typically selects a higher number of inducing points, while OIPS tends to select the fewest, sometimes compromising accuracy. In contrast, our method achieves a balanced selection of inducing points, maintaining comparable accuracy levels to the other methods, with a single parameter setting performing well across all datasets. Figure 3 exemplifies this, showing how our method maintains accuracy with a prudent selection of inducing points. See Appendix F.3 for the rest of the datasets.
Table 1: UCI datasets: Mean (std) over different training/test splits of the number of inducing points (M) and RMSE for the last batch.
Dataset Exact GP Cond. Variance [3] OIPS [14] VIPS (Ours)
RMSE M RMSE M RMSE M RMSE
Concrete .33(.01) 258(85) .4(.03) 50(19) .52(.02) 371(93) .36(.03)
Skillcraft .65(.03) 238(32) .65(.03) 146(25) .65(.03) 139(2) .65(.03)
Kin8nm .27(.00) 5325(22) .27(.00) 1539(183) .29(.01) 2967(31) .27(.01)
Elevators .38(.00) 949(21) .39(.00) 264(0) .39(.01) 329(2) .39(.00)
Bike .20(.00) 4851(42) .22(.00) 366(44) .31(.01) 1021(14) .27(.00)
Figure 3: Plot of the selected number of inducing points and RMSE versus the number of data points observed for different inducing point selection methods using the UCI dataset “kin8nm”. The exact GP which has access to all observations is added as a benchmark for first, middle and last batch.
5.4 Magnetic Anomalies
Finally, we consider the experimental setup described by Chang et al. [4] and the real-world data from Solin et al. [39], where a small wheeled robot with a 3-axis magnetometer navigates a 6m x 6m indoor space to map magnetic field anomalies online. This experiment mimics a real-world setting with an ever-expanding domain, where the robot is not confined to a predefined area. In this context, restricting the model’s capacity is impractical since we need to learn new parts of the space continuously. This requires a method that can accommodate new data without uncontrollably increasing the model’s capacity.
We explore two approaches for learning the robot’s path. In the first approach, each batch represents a complete path taken by the robot. Figure 4(a) shows the final magnetic field estimate after learning from all paths with the marginal variance controlling the opacity. Our method achieves an accurate representation by dynamically adjusting the inducing points. In the second approach, the robot’s path is processed in continuous batches as it moves forward. Figures 4(c) and 4(b) display the mean estimates for the final batch using our method VIPS and Conditional Variance. Our method produces a similar local estimate to Figure 4(a), uses fewer inducing points than Conditional Variance to achieve similar accuracy and does not require hyperparameter tuning. See Appendix F.4 for more details and results.
(a) Learning paths sequentially.
(b) VIPS (Ours).
(c) Conditional Variance
@@ 5.4 Magnetic Anomalies
Figure 4: A small robot is used to perform sequential estimation of magnetic field anomalies. The strength of the magnetic field is given by 10 $\mu\mathrm{T}$ to 90 $\mu\mathrm{T}$. (a)-(b) show the outcome of learning when data for a single path is received continuously. (c) shows the final estimate of the magnitude field learned sequentially through different paths.
@@ 6 Discussion
In this work, we introduce a principled approach for determining the number of inducing points required in streaming sparse GP regression. Unlike conventional methods that rely on a fixed number of inducing points, our approach automatically adjusts the number of inducing points and offers an accuracy guarantee through user-controlled parameters. Our approach effectively eliminates the need to specify a “size” hyperparameter upfront, thereby addressing a significant bottleneck in traditional methods. This allows a GP to be trained without compromising performance due to insufficient memory or wasting computational resources with an unnecessarily large memory allocation. While our current application focuses on GPs we hope this is extendable for the development of adaptive Neural Networks.
Appendix A Societal Impact
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix B Code
The methods discussed in this work, along with the code to reproduce our results, are available online at https://github.com/guiomarpescador/vips.
Appendix C Vegas Inducing Point Selection (VIPS) Algorithm
@@
  Output: Updated set $\mathbf{Z}_{n}=\mathbf{Z}_{o}\cup\{\mathbf{x}_{m^{\prime}}\}_{m^{\prime}=1}^{M_{\mathbf{b}}}$.
@@
  while $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})\geq\delta|\widehat{\mathcal{U}}(M)-\mathcal{L}_{noise}(\hat{\mu},\hat{\sigma})|$ do
@@
Algorithm 1 presents an inducing point selection method using our stopping criterion combined with the location selection strategy “greedy variance” [3, 12, 13]. At each batch, we keep the old inducing point locations fixed and choose the new set of inducing points from among the locations in $\mathbf{X}_{n}$. The method takes as input a value for the hyperparameter $\theta_{n}$. In practice, we will set $\theta_{n}=\theta_{o}$ to select the number of inducing points; the hyperparameter $\theta_{n}$ is inferred by optimising ^$\widehat{\mathcal{L}}(M_{\mathbf{b}})$ once the inducing point locations $\mathbf{Z}_{n}$ has been chosen.
Appendix D Proof of Guarantee

### v4

@@ 5.1 Model size and data distribution
Figure 2: (a) Performance comparison of fixed memory approaches (blue curves with $M=10,20,30$ inducing points) and VIPS, with $M$ (shown at the top) inducing points at each batch. (b) (log) Time taken to train the online GP model on the “naval” dataset divided into 20 batches with fixed size (oracle: $M=100$, heuristic: $M=1000$) and VIPS, our adaptive method.
5.2 The impact of model capacity on accuracy and training time
Here, we compare a fixed to a dynamic model size, with inducing points selected according to the greedy variance criterion (Burt et al., 2020), to demonstrate the challenges of a fixed model size. Since the total dataset size and data characteristics are not known at the start of training, selecting a fixed size will either lead to a model that is too small and underperforms, or a model that is too large and wastes computational resources. Adaptive size solves these issues.
[…]
Training cost comparison: We use the “naval” UCI dataset divided into 20 batches, to compare the training time of models with a fixed and dynamic size. We compare a best-case fixed-size model with the smallest number of inducing points that would still give near-optimal performance. For “naval”, this is $M=100$, which would have to be set by an “oracle”. We also include a typical heuristic, of $M=1000$ (around 1/10th of the total data points), that would ensure sufficient capacity to handle different data patterns and complexities. Finally, we test our adaptive method, VIPS, which adjusts the number of inducing points as needed while achieving near-exact performance for the dataset. As shown in Figure 2(b), VIPS results in lower computational time by only growing its model size as new data is seen, beating even the oracle method.
5.3 Continual learning of UCI datasets
We compare our method, VIPS, to two other inducing point selection approaches: Conditional Variance (CV) and OIPS (Galy-Fajou & Opper, 2021) (see details in App. B). We use six datasets from the UCI repository (Dua & Graff, 2017), simulating a continual learning scenario by sorting the data along the first dimension and dividing it into batches. When evaluating inducing point selection methods, we must consider the trade-off between accuracy and model size: while selecting more inducing points improves model accuracy, it also increases computational cost.
Table 1: Mean (std) of inducing points in final batch across training/test splits at operating point selected to achieve 10% RMSE threshold. Cross (✗) indicates failure to meet accuracy threshold and "Max." denotes reaching maximum capacity limit. Values for OIPS are in italics, as the operating point was selected based on only five datasets, making it less stringent than other methods.
Dataset CV OIPS VIPS (Ours)
Concrete 492(54) 240(102) 234(116)
Skillcraft 739(34) 174(16) 134(9)
Kin8nm 6316(10) 6458(8) 1904(41)
Naval 49(1) 57(1)
Elevators 3133(140) 298(8) 291(4)
Bike Max. 7000 1964(137) 650(17)
Table 2: Mean (std) of inducing points in final batch across training/test splits at operating point selected to achieve 10% RMSE threshold in all but one dataset. ✗ Gray shows cases exceeding threshold, with percentages showing % difference to full-batch GP relative to a noise model.
Dataset CV OIPS VIPS (Ours)
Concrete 184(74) 240(102) 178(95)
Skillcraft 226(22) 174(16) 133(0)
Kin8nm 5334(27) 6458(8) 1563(39)
Naval ✗ 23(8)-323% ✗ 16(1)-315.28% 48(3)
Elevators 836(27) 298(8) 281(4)
Bike 4895(28) 1964(137) ✗  549(14)-10.8%
Table 3: Mean (std) of inducing points in final batch across training/test splits at operating point selected to achieve 10% RMSE threshold. Cross (✗) indicates failure to meet accuracy threshold and "Max." denotes reaching maximum capacity limit. Values for OIPS are in italics, as the operating point was selected based on only five datasets, making it less stringent than other methods.
Dataset CV OIPS VIPS (Ours)
Concrete 492(54) 383(90) 451(77)
Skillcraft 1091(107) 236(38) 195(18)
Kin8nm 6469(7) 6523(9) 5065(46)
Naval 903(8) 509(15)
Elevators 3133(140) 386(47) 487(24)
Bike Max. 7000 3916(122) 2794(148)
Table 4: Mean (std) of inducing points in final batch across training/test splits at operating point selected to achieve 10% NLPD threshold in all but one dataset. ✗ Gray shows cases exceeding threshold, with percentages showing % difference to full-batch GP relative to a noise model.
Dataset CV OIPS VIPS (Ours)
Concrete 303(78) 383(90) 434(79)
Skillcraft 295(13) 236(38) 177(13)
Kin8nm 5685(26) 6523(9) 4281(69)
Naval ✗ 27(6)-2406% ✗ 20(1)-133.87% ✗  373(19)-11.92%
Elevators 1099(50) 386(47) 438(17)
Bike 5563(36) 3916(122) 2296(138)
Our goal is to minimise computational cost, while satisfying the constraint that we reach a fixed target accuracy threshold (e.g. within 10% RMSE of a full-batch GP). To benchmark our methods, we select a single hyperparameter across datasets, so that our accuracy constraint is always met, while otherwise using the fewest inducing points. We refer to this hyperparameter as the operating point. For this operating point, we then measure the number of inducing points that the method chooses for each dataset, with a lower result being better. If across datasets, there is a large variation in hyperparameter values needed to achieve the accuracy constraint, selecting the most stringent criterion will cause large numbers of inducing points to be used for other datasets. This simulates a realistic setting, where one cannot know what the optimal hyperparameter value is for each dataset beforehand, and penalises methods that are hard to tune, due to large variations in the required hyperparameter values across datasets.
Table 1 shows the number of inducing points for each method at the operating point that achieves the accuracy threshold within 10% RMSE of a full-batch GP. VIPS requires fewer inducing points in the majority of datasets. CV often results in larger models, reaching the 7000-point limit on Bike. OIPS performs inconsistently, matching VIPS on smaller datasets but oversizing on larger ones. For the Naval dataset, OIPS failed to meet the accuracy constraint across all tested hyperparameter values. While extending the hyperparameter range might have improved accuracy, it would have led to even larger models. Therefore, we selected an operating point that satisfied our conditions for five out of six datasets, treating Naval as an outlier. Table 2 shows the results when this exception is applied to all methods. We observe a similar trend, with VIPS achieving the smallest model sizes. Tables 3 and 4 report analogous results for the threshold within 10% NLPD of a full-batch GP. The same pattern holds. Figure 3 summarises these outcomes across outlier allowances, with VIPS more frequently achieving minimal model size. App. D.3 provides additional results under different thresholds.
Our results show that VIPS performs consistently across datasets, allowing a single pre-set hyperparameter to be used. In contrast, OIPS and CV require tuning, as no single hyperparameter works well for all datasets. This is a limitation in continual learning, where the whole dataset is unavailable in advance.
Figure 3: Number of datasets where each method achieves minimal model size at different outlier allowances. A “win” is assigned to the dataset with the smallest model which satisfies the (a) RMSE and (b) NLPD thresholds of 10%. Counts represent absolute wins. Higher counts indicate better method robustness across datasets.
5.4 Continual Learning of Magnetic Field
We apply the optimal hyperparameters found in the previous section, selected under the 10% RMSE threshold, to the experimental setup from Chang et al. (2023a) using real-world data from Solin et al. (2018), where a robot maps magnetic field anomalies in a 6m x 6m indoor space. This scenario simulates an expanding domain, where the model needs to continuously learn new areas without uncontrollably increasing its capacity. Our goal is to test whether the methods can form representations by progressively spreading inducing points and learning hyperparameters without prior tuning in two scenarios: 1) sequentially mapping multiple robot trajectories and 2) processing a single path in continuous batches. Figure 4 shows the final magnetic field estimates produced by VIPS in both experimental settings. Comparative results for CV and OIPS are provided in App. D.4. The results show that VIPS was able to grow its memory as the robot moved, allowing for accurate modelling of the path and increasing its capacity dynamically. In contrast, we observe that CV increases its capacity uncontrollably while OIPS provides a less accurate estimate of the magnetic field due to choosing fewer inducing points.
(a) Learning paths 1, 2, 4, 5 sequentially with VIPS.
(b) Learning path 3 with VIPS. M: 134, RMSE: 7.55.
@@ 5.4 Continual Learning of Magnetic Field
Figure 4: A small robot is used to perform sequential estimation of magnetic field anomalies. The strength of the magnetic field is given by 10 $\mu\mathrm{T}$ to 90 $\mu\mathrm{T}$. (a) shows the final estimate of the magnitude field learned sequentially through different paths. (b) show the outcome of learning a single path continuously with the black dots representing the chosen inducing points.
@@ 6 Discussion
In this work, we propose a method to dynamically adjust the number of inducing variables in streaming GP regression. Our method achieves a performance close to full-batch approaches while minimising model size. It relies on a single hyperparameter to balance accuracy and complexity, and we demonstrate that our method performs consistently across diverse datasets. This reduces the need for extensive hyperparameter tuning and eliminates the requirement to pre-define model size, thereby addressing a significant bottleneck in traditional methods. While our current application focuses on GPs we hope this is extendable for the development of adaptive neural networks.
Acknowledgements
GPB is supported by EPSRC through the Statistical Machine Learning (StatML) CDT programme, grant no. EP/S023151/1.
Impact Statement
@@ Acknowledgements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Code
The methods discussed in this work, along with the code to reproduce our results, are available online at https://github.com/guiomarpescador/vips. All experiments and figures were done in Python using libraries TensorFlow v.2.12 (Abadi et al., 2016) and GPflow v.2.9 (Matthews et al., 2017). Both packages are under Apache License 2.0. The implementation of the online lower bound (see eq. (8)) by Bui et al. (2017) is available on GitHub under Apache license 2.0. For the “greedy variance” selection criterion and algorithm, we use the implementation provided in Burt et al. (2020) available under Apache license 2.0.
Appendix B Adaptive Inducing Points Selection Methods
In sections 5.3 and 5.4 we compare our method, VIPS, to two other adaptive approaches: Conditional Variance (CV) and OIPS (Galy-Fajou & Opper, 2021). The key distinction between these approaches lies in their selection strategy of inducing points. OIPS uses a unified selection process (Algorithm 2), where the choice of both location and number of inducing points is determined by a single metric. In particular, each point $\mathbf{x}_{i}$ in the training set $X_{n}$ is evaluated individually, and those that meet the threshold condition are added to the inducing set $\mathbf{Z}_{n}$. In contrast, both CV and VIPS use a decoupled selection process (Algorithm 3), which separates two decisions: where to place inducing points and how many to include. These methods iteratively select the best candidate for the next inducing point from a pool of potential locations, continuing until a stopping criterion is met, allowing for more control over the model size. VIPS and CV share the same greedy variance selection criterion but differ in their candidate pools and stopping criteria. Table 5 summarises the properties of these methods, including the fixed memory approach of Bui et al. (2017) for comparison. The next three subsections provide detailed descriptions of each method’s specific implementation.
Algorithm 2 Unified inducing set selection process done at each batch.
  Input: New batch of data $\mathbf{X}_{n}$, current set inducing points $\mathbf{Z}_{old}$.
  Output: Updated set $\mathbf{Z}_{n}$
  Initialise $\mathbf{Z}_{n}=\mathbf{Z}_{old}$
  for $\mathbf{x}_{i}\in\mathbf{X}_{n}$ do
     if addition_criterion($\mathbf{x}_{i}$) < threshold then
        $\mathbf{Z}_{n}=\mathbf{Z}_{n}\cup\{\mathbf{x}_{i}\}$
     end if
  end for
Algorithm 3 Decoupled inducing set selection process done at each batch.
  Input: New batch of data $\mathbf{X}_{n}$, current set inducing points $\mathbf{Z}_{old}$.
  Output: Updated set $\mathbf{Z}_{n}$
  Set candidate pool $\mathbf{X}_{pool}$ based on method (see Table 5) and initialise $\mathbf{Z}_{n}$ accordingly.
  while stopping_criterion($\mathbf{Z}_{n}$) $>$ threshold do
     $\mathbf{x}_{best}=$ selection_criterion($\mathbf{X}_{pool}$, $\mathbf{Z}$) {Greedy variance criterion}
     $\mathbf{Z}_{n}=\mathbf{Z}_{n}\cup\{\mathbf{x}_{best}\}$
  end while
Method Model Size Selection Pool Selection Criterion Stopping Criterion
Bui et al. (2017) Fixed $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ Gradient optimisation M constant
OIPS (Galy-Fajou & Opper, 2021) Adaptive $\{\mathbf{X}_{n}\}$ $\max\mathbf{k}_{\mathbf{u}}(\mathbf{x})\leq\rho$ $\max\mathbf{k}_{\mathbf{u}}(\mathbf{x})\leq\bar{\rho}$
Cond. Variance (CV) Adaptive $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ Greedy variance $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)\leq\eta$.
VIPS Adaptive $\{\mathbf{X}_{n}\}$ Greedy variance $\widehat{\mathcal{U}}-\widehat{\mathcal{L}}\leq\alpha$
B.1 Vegas Inducing Point Selection (VIPS) Algorithm
To select the location of our new inducing points we use the location selection strategy “greedy variance” proposed in Burt et al. (2020). This strategy iteratively selects points from a set based on a preference criterion until a stopping condition is met. In particular, it chooses the location of the next inducing point to maximise the marginal variance in the conditional prior $p(f_{\neq\mathbf{u}}|\mathbf{u})$. This is equivalent to maximising $\operatorname{diag}[\mathbf{K_{ff}}-\mathbf{Q_{ff}}]$. In continual learning, Chang et al. (2023a) use the “greedy variance” criterion by defining $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ as the selection pool from which inducing point locations are selected and maintaining a fixed number of inducing points. Similarly, Maddox et al. (2021) extends the “greedy variance” criterion to heteroskedastic Gaussian likelihoods and also uses a fixed memory approach. In our case, we tested the location strategy with our stopping criterion using both $\{\mathbf{Z}_{o},\mathbf{X}_{n}\}$ and $\{\mathbf{X}_{n}\}$ as candidates pool for the locations of the inducing points. We did not find a substantial difference between the methods and hence opted for the simpler version where we keep the old inducing point locations fixed and choose the new set of inducing points from among the locations in $\{\mathbf{X}_{n}\}$.
Algorithm 4 presents an inducing point selection method using our stopping criterion combined with the location selection strategy “greedy variance”. The method takes as input a value for the hyperparameter $\theta_{n}$. In practice, we will set $\theta_{n}=\theta_{o}$ to select the number of inducing points; the hyperparameter $\theta_{n}$ is inferred by optimising ^$\widehat{\mathcal{L}}(M_{\mathbf{b}})$ once the inducing point locations $\mathbf{Z}_{n}$ has been chosen. In Algorithm 1, $\widehat{\mathcal{U}}(M)$ is used to calculate the stopping criterion. However, in practice, since for the continual learning settings we investigate $N_{n}$ is small enough, we will use $\widehat{\mathcal{U}}(M_{\mathbf{a}}+N_{n})=\mathcal{L}^{*}$. This value is calculated once at the beginning of the process. The algorithm’s complexity depends on the number of inducing points $M_{\mathbf{b}}$ used to compute the lower bound ^$\widehat{\mathcal{L}}(M_{\mathbf{b}})$ at each iteration. The computational complexity for calculating $\widehat{\mathcal{L}}$ at each batch is $O(N_{n}M_{\mathbf{b}}^{2}+M_{\mathbf{b}}^{3})$, and the memory requirement is $O(M_{\mathbf{b}}^{2})$, where $M_{\mathbf{b}}$ represents the total number of inducing points in the current batch.
@@ B.1 Vegas Inducing Point Selection (VIPS) Algorithm
  Output: Updated set $\mathbf{Z}_{n}=\mathbf{Z}_{o}\cup\{\mathbf{x}_{m^{\prime}}\}_{m^{\prime}=1}^{M^{\prime}}$, where $|\mathbf{Z}_{n}|=M_{\mathbf{b}}$.
@@
  while $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})\leq\delta|\widehat{\mathcal{U}}(M)-\mathcal{L}_{noise}(\hat{\mu},\hat{\sigma})|$ do
@@
B.2 Conditional Variance
The implementation of the Conditional Variance method is presented in Algorithm 5. This method uses the “greedy variance” strategy that iteratively chooses the location of the next inducing point. As a stopping criterion, it uses the quantity $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)$. In this algorithm, new inducing points are no longer added once $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)$ falls below a chosen tolerance value $\eta$. Although, this approach was mentioned in Burt et al. (2020), this stopping criterion has not yet been tested in the literature. The hyperparameter $\eta$ is determined by the user.
Algorithm 5 Conditional Variance (CV)
  Input: $\mathbf{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{N_{n}}$, $\mathbf{Z}_{o}=\{\mathbf{z}_{m}\}_{m=1}^{M_{\mathbf{a}}}$, $\theta_{n}$, kernel, $k(\cdot,\cdot|\theta_{n})$, threshold $\eta$.
  Output: Updated set of inducing points $\mathbf{Z}_{n}=\{\mathbf{z}_{m}\}_{m=1}^{M^{\prime}}\cup\{\mathbf{x}_{m^{\prime}}\}_{m^{\prime}=1}^{M_{b}-M^{\prime}}$, where $|\mathbf{Z}_{n}|=M_{\mathbf{b}}$.
  Initialise location selection pool: $\mathbf{X}_{pool}=\mathbf{Z}_{o}\cup\mathbf{X}_{m}$.
  Initialise $\mathbf{Z}_{n}={\operatorname{argmax}}_{\mathbf{x}\in\mathbf{X}_{pool}}\,k(\mathbf{x},\mathbf{x})$.
  while $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)\leq\eta$ do
     Select $\mathbf{x}={\operatorname{argmax}}_{\mathbf{x}\in\mathbf{X}_{pool}}\,k(\mathbf{x},\mathbf{x})-\mathbf{k}_{\mathbf{b}}(\mathbf{x})^{\top}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{k}_{\mathbf{b}}(\mathbf{x})$.
  end while
B.3 Online Inducing Point Selection (OIPS)
Galy-Fajou & Opper (2021) introduced the Online Inducing Points Selection (OIPS) algorithm, which iteratively adds points from $\mathbf{X}_{n}$ to the set of inducing points. The algorithm assesses the impact of each new point on the existing inducing set, based on a covariance threshold. A point $\mathbf{x}$ is added if the maximum value of $\mathbf{k}_{\mathbf{u}}(\mathbf{x})$ falls below a user-defined threshold $\rho$. While the original algorithm implicitly assumes unit kernel variance, Algorithm 6 presents our implementation where we scale the threshold $\rho$ by the kernel variance $\sigma^{2}_{f}$ to $\bar{\rho}=\sigma^{2}_{f}\rho$. This modification ensures the threshold comparison remains valid for arbitrary kernel variances.
Algorithm 6 Online Inducing Point Selection (OIPS)
  Input: $\mathbf{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{N_{n}}$, $\mathbf{Z}_{o}=\{\mathbf{z}_{m}\}_{m=1}^{M_{\mathbf{a}}}$, kernel function $k(\cdot,\cdot|\theta_{n})$, kernel hyperparameters $\theta_{n}$ (including variance $\sigma_{f}^{2}$), acceptance threshold $10<\rho<$ 1.
  Output: Updated set $\mathbf{Z}_{n}=\mathbf{Z}_{o}\cup\{\mathbf{x}_{m^{\prime}}\}_{m^{\prime}=1}^{M^{\prime}}$, where $|\mathbf{Z}_{n}|=M_{\mathbf{b}}$.
  Initialise $\mathbf{Z}_{n}=\mathbf{Z}_{o}$.
  Initialise $\bar{\rho}=\rho\cdot\sigma^{2}_{f}$.
  for all $\mathbf{x}_{i}\in\mathbf{X}_{n}$ do
     $d=\max_{j}\left(k\left(\mathbf{x}_{i},\mathbf{z}_{j}|\theta_{n}\right)\right),\,\forall\,\mathbf{z}_{j}\in\mathbf{Z}_{n}$.
     if $d<\bar{\rho}$ then
        Add $\mathbf{x}_{i}$ to the set of inducing points: $\mathbf{Z}_{n}=\mathbf{Z}_{n}\cup\{\mathbf{x}_{i}\}$.
     end if
  end for
Appendix C Proof of Guarantee